OpenStudy (anonymous):

How do you solve for x: x-2(x+10)=12 I tried distributive property and combining like terms but I don't know what to do.

OpenStudy (anonymous):

First multiply x-2x-20=12 because infront of 2 there's a (-) sign so you need to distribute that as well. Now x-2x=-x, so you write, -x-20=12 When you take 20 on the other side, you get -x=12+20=32 Multiply by (-1) and you get x=-32. Proof: -32-2(-32+10)=12 -32-2(-22)=12 -32+44=12 12=12

OpenStudy (anonymous):

Thank you! that really helped. I realized that what I did was I forgot when I was doing the distributive property, to multiply 2(10). Thanks again

OpenStudy (anonymous):

Hehe, dont worry. Practice makes perfect!